3. A regular square pyramid has a height of 8m. If the slant height makes an angle of 45º with the base, find the lateral area of the pyramid.
1
Expert's answer
2020-12-08T19:18:07-0500

Consider a regular square pyramid

Let "a=" the base edge, "h=" the altitude, and "s=" the slant height.

The lateral area "A_L" of a regular square pyramid is


"A_L=\dfrac{1}{2}(4a)s"


The slant height makes an angle "\alpha=45\degree" with the base.

From the right triangle


"\sin\alpha=\dfrac{h}{s}, \tan \alpha=\dfrac{h}{a/2}"


"s=\dfrac{h}{\sin\alpha}=\dfrac{8}{\sin45\degree}=8\sqrt{2}(m)"

"a=\dfrac{2h}{\tan\alpha}=\dfrac{2(8)}{\tan45\degree }=16(m)"


"A_L=2as=2(16)(8\sqrt{2})=256\sqrt{2}(m^2)"

"A_L=256\sqrt{2}m^2"